Your lab manager asks you to create the new control range for the new lot of hematology control level 1.  The following 20 hemoglobin values are gathered to determine the control limits.  Using a calculator or website SD function, calculate and record the mean, the standard deviation, and the 95% confidence interval for this set of values.





4.2

4.7

4.3

4.4

4.6

4.7

4.9

5.0

4.6

4.0

4.4

4.5

4.0

4.9
4.7

4.1

4.2

4.7

     4.8

     4.4





Mean:__________   1 SD:__________      95% Confidence range:________________


Given the results above, label the following Levey-Jennings Control chart with all of the following:

Mean

Values for -3SD, -2SD, -1SD, +1SD, +2SD, +3SD

.


You have your Levey-Jennings Control chart ready.  Your supervisor tells you to begin using the lot.  Record the first 10 days of Control 1 Hemoglobin results.  In your lab, you run controls once per day because you are a clinic open from 8-5pm.  Plot each data point on your labeled chart and circle THOSE controls THAT ARE considered “out.”






                                   DAY of August

1= 4.5 mg/dl

2=4.2 mg/dl

3=4.1 mg/dl

4=4.6 mg/dl

5=4.7 mg/dl


6=4.8 mg/dl

7=4.4 mg/dl

8=4.4 mg/dl

9=4.2 mg/dl

10=4.9 mg/dl

ACCEPTABLE                    OUT OF CONTROL

You run a normal serum glucose control and obtain a value of 78.4 mg/dL.  The mean for this control is 74.8 mg/dL and 1 SD has been calculated as 1.2.  Is the value of 78.4 within 2SD (95% confidence interval)?  Circle the correct answer.

YES         NO







  







PART TWO

A hematology laboratory is using data reported from 2 different analyzers to determine which is more accurate for the testing of samples in their hospital.  The following data is compiled on 30 runs of normal control.  The analytes being compared are the WBC, RBC, HGB, HCT and platelet values.

Using the EXCEL program you were introduced in the computer lab; calculate the mean, standard deviation and coefficient of variation for each analyte.  Which analyzer is more accurate?  What led you to this conclusion?  Plot you data using the Levy-Jennings chart.  Note any shifts or trends.

Analyzer #1

WBC        RBC        HGB        HCT        Platelet

7.9        4.23        12.4        36.4        254

7.8        4.25        12.3        36.6        244

7.7        4.25        12.4        36.5        238

7.8        4.22        12.3        36.6        249

7.7        4.26        12.2        36.3        244

7.6        4.23        12.4        36.5        230

7.4        4.23        12.5        36.3        251

7.9        4.21        12.4        36.4        253

7.8        4.19        12.6        36.9        243

7.9        4.18        12.7        36.9        241

8.0        4.18        12.5        36.3        239

7.8        4.16        12.2        35.9        238

7.8        4.15        12.1        35.7        236

7.9        4.22        12.4        36.7        231

8.0        4.26        12.4        36.3        246

7.9        4.23        12.4        36.4        253

7.8        4.18        12.4        36.1        258

7.9        4.13        12.4        35.6        261

7.9        4.20        12.4        36.3        255

8.1        4.26        12.6        36.6        264

7.7        4.18        12.3        36.2        233

7.8        4.24        12.3        36.9        233

8.0        4.20        12.3        36.5        246

7.9        4.20        12.3        36.3        242

7.8        4.23        12.3        36.4        240

7.9        4.22        12.3        36.3        243

7.7        4.26        12.5        36.9        238

7.8        4.15        12.2        35.8        241

7.9        4.23        12.3        36.6        234

7.9        4.20        12.4        36.2        253

Analyzer #2

WBC        RBC        HGB        HCT        Platelet

7.8        4.23        12.5        36.4        247

7.9        4.17        12.4        36.0        250

7.8        4.23        12.3        36.5        251

7.9        4.19        12.3        36.0        259

8.1        4.23        12.4        36.1        258

8.1        4.23        12.3        36.5        256

7.8        4.21        12.4        36.4        260

7.9        4.21        12.3        36.3        263

7.9        4.21        12.5        36.2        266

8.0        4.21        12.5        36.3        265

7.8        4.23        12.1        36.6        248

7.8        4.19        12.1        36.2        233

7.8        4.19        12.4        36.4        248

7.6        4.15        12.2        35.9        241

7.7        4.21        12.3        36.6        251

7.8        4.16        12.2        36.4        233

7.8        4.20        12.3        36.5        256

7.8        4.24        12.3        36.7        244

7.8        4.19        12.4        36.1        245

8.0        4.26        12.5        37.0        251

7.9        4.21        12.4        36.6        250

7.7        4.19        12.3        36.3        254

7.8        4.22        12.3        36.6        250

7.7        4.19        12.3        36.0        252

7.8        4.16        12.2        35.9        244

7.9        4.16        12.3        35.8        260

7.9        4.20        12.3        36.3        251

7.7        4.18        12.3        36.0        251

7.8        4.23        12.4        36.5        258

7.6        4.21        12.5        36.7        256


Your lab manager asks you to create the new control range for the new lot of hematology control level 1.  The following 20 hemoglobin values are gathered to determine the control limits.  Using a calculator or website SD function, calculate and record the mean, the standard deviation, and the 95% confidence interval for this set of values.





4.2

4.7

4.3

4.4

4.6

4.7

4.9

5.0

4.6

4.0

4.4

4.5

4.0

4.9
4.7

4.1

4.2

4.7

     4.8

     4.4





Mean:__________   1 SD:__________      95% Confidence range:________________


Given the results above, label the following Levey-Jennings Control chart with all of the following:

Mean

Values for -3SD, -2SD, -1SD, +1SD, +2SD, +3SD

.


You have your Levey-Jennings Control chart ready.  Your supervisor tells you to begin using the lot.  Record the first 10 days of Control 1 Hemoglobin results.  In your lab, you run controls once per day because you are a clinic open from 8-5pm.  Plot each data point on your labeled chart and circle THOSE controls THAT ARE considered “out.”






                                   DAY of August

1= 4.5 mg/dl

2=4.2 mg/dl

3=4.1 mg/dl

4=4.6 mg/dl

5=4.7 mg/dl


6=4.8 mg/dl

7=4.4 mg/dl

8=4.4 mg/dl

9=4.2 mg/dl

10=4.9 mg/dl

ACCEPTABLE                    OUT OF CONTROL

You run a normal serum glucose control and obtain a value of 78.4 mg/dL.  The mean for this control is 74.8 mg/dL and 1 SD has been calculated as 1.2.  Is the value of 78.4 within 2SD (95% confidence interval)?  Circle the correct answer.

YES         NO







  







PART TWO

A hematology laboratory is using data reported from 2 different analyzers to determine which is more accurate for the testing of samples in their hospital.  The following data is compiled on 30 runs of normal control.  The analytes being compared are the WBC, RBC, HGB, HCT and platelet values.

Using the EXCEL program you were introduced in the computer lab; calculate the mean, standard deviation and coefficient of variation for each analyte.  Which analyzer is more accurate?  What led you to this conclusion?  Plot you data using the Levy-Jennings chart.  Note any shifts or trends.

Analyzer #1

WBC        RBC        HGB        HCT        Platelet

7.9        4.23        12.4        36.4        254

7.8        4.25        12.3        36.6        244

7.7        4.25        12.4        36.5        238

7.8        4.22        12.3        36.6        249

7.7        4.26        12.2        36.3        244

7.6        4.23        12.4        36.5        230

7.4        4.23        12.5        36.3        251

7.9        4.21        12.4        36.4        253

7.8        4.19        12.6        36.9        243

7.9        4.18        12.7        36.9        241

8.0        4.18        12.5        36.3        239

7.8        4.16        12.2        35.9        238

7.8        4.15        12.1        35.7        236

7.9        4.22        12.4        36.7        231

8.0        4.26        12.4        36.3        246

7.9        4.23        12.4        36.4        253

7.8        4.18        12.4        36.1        258

7.9        4.13        12.4        35.6        261

7.9        4.20        12.4        36.3        255

8.1        4.26        12.6        36.6        264

7.7        4.18        12.3        36.2        233

7.8        4.24        12.3        36.9        233

8.0        4.20        12.3        36.5        246

7.9        4.20        12.3        36.3        242

7.8        4.23        12.3        36.4        240

7.9        4.22        12.3        36.3        243

7.7        4.26        12.5        36.9        238

7.8        4.15        12.2        35.8        241

7.9        4.23        12.3        36.6        234

7.9        4.20        12.4        36.2        253

Analyzer #2

WBC        RBC        HGB        HCT        Platelet

7.8        4.23        12.5        36.4        247

7.9        4.17        12.4        36.0        250

7.8        4.23        12.3        36.5        251

7.9        4.19        12.3        36.0        259

8.1        4.23        12.4        36.1        258

8.1        4.23        12.3        36.5        256

7.8        4.21        12.4        36.4        260

7.9        4.21        12.3        36.3        263

7.9        4.21        12.5        36.2        266

8.0        4.21        12.5        36.3        265

7.8        4.23        12.1        36.6        248

7.8        4.19        12.1        36.2        233

7.8        4.19        12.4        36.4        248

7.6        4.15        12.2        35.9        241

7.7        4.21        12.3        36.6        251

7.8        4.16        12.2        36.4        233

7.8        4.20        12.3        36.5        256

7.8        4.24        12.3        36.7        244

7.8        4.19        12.4        36.1        245

8.0        4.26        12.5        37.0        251

7.9        4.21        12.4        36.6        250

7.7        4.19        12.3        36.3        254

7.8        4.22        12.3        36.6        250

7.7        4.19        12.3        36.0        252

7.8        4.16        12.2        35.9        244

7.9        4.16        12.3        35.8        260

7.9        4.20        12.3        36.3        251

7.7        4.18        12.3        36.0        251

7.8        4.23        12.4        36.5        258

7.6        4.21        12.5        36.7        256




watch Jening QC Charts and from there it would help resolve the problem I guess
Answer #1

Mean value of Hemoglobin = 412+ 4.7+ 4.3+ 4.444.6+4 7+4 9+5.0+ 4.6+4.ot 4,4+ 4,5 +4.0+4.9+4.7+4.14 4.2+4.7+ 4.8+4,4 = 90.1 Me(X13-4)² =(4,0-4,5)2 = 0,25 (X14-uy ² = (4.9-4.5; 2 = 0.16 (x15-4)2 = (4.7-4.5 2 = 0.04 (X16-4² = (4,1-4,5)2 = 0.16 (X17-4)2so Mean = 4.5 ISD = 0.29 95% confidence range = ± 0.127 so #1ST TISD = 4.79 +2SD = 5.08 +38D = 5.37 - ISD = 4. 21 - 2 SD = 3.